# Toward a 2D Local Implementation of Quantum LDPC Codes

### v1

@@ IV.4 Circuit-level simulations
Fig. 7(a)–(b) shows the results of these simulations for several codes listed in Table 1. As a comparison, we perform the same simulations with the rotated surface code which has parameters $[[d^{2},1,d]]$. To decode, we follow the same process as described in Section IV.3 but instead use the minimum-weight perfect matching decoder [78]. As the BB codes encode multiple logical qubits in a single block, multiple copies of the surface code must be used to achieve the same number of logical qubits. If $p_{SC,1}$ is the logical error rate of simulating a single rotated surface code for $t$ error correction rounds, then $k$ copies of the surface code have a logical error rate
@@
In addition to the logical error rate, another important performance metric is the number of qubits used to achieve it. For the BB codes and the bilayer architecture, this includes the ancillary check qubits as well as the entire routing layer, which for an $[[n,k,d]]$ code uses $4n$ qubits in total. The rotated surface code uses $d^{2}-1$ additional check qubits, which brings the total number of qubits to $12d^{2}-$ 1 for each copy. The total number of qubits used is listed together with the code parameters in Fig. 7. The error bars on the data points are calculated using the standard error when sampling from a binomial distribution $\sqrt{p_{\log}(1-p_{\log})/N}$, where $N$ is the number of collected samples. Additionally, we plot a fit of
@@
The smallest BB codes encoding $k=8$ logical qubits are outperformed by surface codes that use fewer physical qubits. However, increasing the block length yields BB codes that approach the performance of similarly sized surface codes. In panel Fig. 7(b), we increase the number of logical qubits to $k=12$ and see advantages to using the proposed architecture in terms of the number of physical qubits used. Compared to twelve patches of a $[[36,1,6]]$ rotated surface code using a total of 852 physical qubits and a logical error rate per round of $\epsilon_{L}=1.43\times 10^{-4}$, we find a $[[144,12,12]]$ BB code using 576 qubits that matches the performance, with $\epsilon_{L}=1.56\times 10^{-4}$. Additionally, we find a $[[196,12,8]]$ code using 784 qubits that outperforms it with $\epsilon_{L}=7.89\times 10^{-5}$. At this scale, the improvements are not so drastic, but we expect to see greater overhead benefits as the block length and number of logical qubits increase.
In Fig. 7(c), we vary the interval at which the long-range generators are measured. For the $[[90,8,6]]$ code, the 35 short-range generators of a single type can be routed, purified, and measured in 9 steps; whereas it takes 5 steps to route, purify, and measure the remaining 10 long-range generators of the same type. This means that measuring the long-range generators every error correction round requires a circuit depth that is 28.5% longer than if the long-range generators were measured every five rounds, at a negligible increase in the logical error rate per round from $\epsilon_{L}=8.41\times 10^{-4}$ to $8.92\times 10^{-4}$. Increasing the size of the code to $[[196,12,8]]$, we again see negligible differences in logical error per round performance between the two measurement schedules, from $\epsilon_{L}=7.41\times 10^{-5}$ to $7.89\times 10^{-5}$, with the additional benefit of a 32.0% decrease in the depth of the syndrome extraction circuit when measured every five rounds. As the block length increases, so does the discrepancy between the measurement times of the short- and long-range generators, increasing the circuit depth savings. Additionally, this discrepancy would disproportionally introduce more errors during the long-range measurement rounds, potentially making it more performant to measure these large generators even less frequently.
@@
However, achieving more significant reductions in circuit depth requires measuring the long-range generators much less frequently, as shown in Fig. 8. We display the potential circuit depth savings for two BB code instances as a function of how many error correction rounds elapse between measurements of the long-range generators. The horizontal red lines indicate the maximum potential savings, corresponding to a schedule where the long-range generators are only measured once at the end of the circuit. For example, the $[[144,12,12]]$ BB code requires 16 (15) steps to measure the short- (long)-range generators of a single type, which gives a maximum depth savings of 48.4%. When measuring every five rounds, as in Fig. 7(a)–(b) and Fig. 9, we see a circuit depth savings of 38.7%, indicated by the vertical gray line. Measuring the long-range generators very infrequently will significantly degrade the error correction performance and may not be worth the reduced circuit depth. Instead, it may be more advantageous to measure the long-range generators relatively frequently, e.g., every 2–5 rounds; in that regime we still see considerable circuit depth savings (50% to 80% of the theoretical maximum), but the impact on the logical error rate is negligible.
@@
Even with the reduced idle error rate that we consider here, idle errors are a significant error source, especially on rounds where the long-range generators are measured. Further decreasing this error rate would substantially reduce the effect of these errors. In Fig. 9, we perform the same simulations as described above, but do not apply idling errors. With idle errors, the $k=8$ BB codes were outperformed by surface codes, whereas without idle errors, we find that all instances achieve comparable logical error rates with fewer physical qubits. Indeed, the $[[150,8,8]]$ code using 600 physical qubits sees an $\times8$ improvement in the logical error rate per round, from $\epsilon_{L}=5.31\times 10^{-5}$ to $5.98\times 10^{-6}$, and now outperforms eight patches of a $[[64,1,8]]$ rotated surface code using $1016$ physical qubits with $\epsilon_{L}=7.51\times 10^{-6}.$ Achieving negligible idle error rates may not be feasible, but it illustrates the regime where our protocol performs best.
@@ V Discussion
We do find BB codes where the same parity check structure is shared between codes of increasing block length, so the time to route the generators is a constant as the code size increases. However, this also means that the percentage of long-range generators and, by extension, the amount of nonlocality in the code, decreases. References [13, 14] showed that it is impossible to beat the asymptotic scaling of the surface code parameters without increasing the amount of nonlocality. Thus to have families of codes with a favorable scaling of $k$ and $d$, the syndrome extraction time in this model must also be increasing.
@@
If the aforementioned issues can be solved, then scaling up should increase the advantage of qLDPC codes over the surface code. One potential solution is to improve the circuit depth of the protocol. An architectural feature that could accomplish this is the ability to perform two-qubit gates on qubits that are some distance $R$ apart [19]. This is a natural operation on neutral-atom devices, where Rydberg-Rydberg interactions, especially dipolar ones [83], can be quite long-range. Such a feature could also be feasible in superconducting devices through the use of medium-range couplers or photonic interconnects [84]. When $R$ is a constant, the asymptotic behavior will remain unchanged; however, practically this would mean that the short-range generators would be much easier to implement. With an appropriate choice of $R$, it would then be possible to use the depth-7 circuit of Ref. [18] to measure the short-range generators, in which case the only difficulty would be to measure the long-range generators in the proposed manner. An alternative approach would be to add additional ancilla layers to the architecture. Although this would further increase the qubit overhead, it would allow for more parallelization during the syndrome measurement, decreasing the total circuit depth.

### v2

@@ IV.4 Circuit-level simulations
Table 2 and Fig. 11(a)–(b) shows the results of these simulations for several codes listed in Table 1. As a comparison, we perform the same simulations with the rotated surface code which has parameters $[[d^{2},1,d]]$. To decode, we follow the same process as described in Section IV.3 but instead use the minimum-weight perfect matching decoder [84]. As the BB codes encode multiple logical qubits in a single block, multiple copies of the surface code must be used to achieve the same number of logical qubits. If $p_{SC,1}$ is the logical error rate of simulating a single rotated surface code for $t$ error correction rounds, then $k$ copies of the surface code have a logical error rate
@@
In addition to the logical error rate, another important performance metric is the number of qubits used to achieve it. For the BB codes and the bilayer architecture, this includes the ancillary check qubits as well as the entire routing layer, which for an $[[n,k,d]]$ code uses $4n$ qubits in total. The rotated surface code uses $d^{2}-1$ additional check qubits, which brings the total number of qubits to $12d^{2}-$ 1 for each copy. The total number of qubits used is listed together with the code parameters in Fig. 11. The error bars on the data points are calculated using the standard error when sampling from a binomial distribution $\sqrt{p_{\log}(1-p_{\log})/N}$, where $N$ is the number of collected samples. Due to the large number of shots taken, $N\sim 10^{5}$, error bars in Fig. 11 and Fig. 12 are nearly invisible. Additionally, we plot a fit of
@@
The smallest BB codes encoding $k=8$ logical qubits are outperformed by surface codes which use fewer physical qubits. However, increasing the block length yields BB codes that surpass the performance of similarly sized surface codes. This is illustrated in Fig. 7, where we see the BB codes achieving a lower logical error rate per round than the surface code while utilizing fewer qubits. Increasing the number of logical qubits to $k=12$, BB codes and the proposed architecture immediately outperform the surface codes in terms of logical error rate and space overhead. Compared to twelve patches of a $[[36,1,6]]$ rotated surface code using a total of 852 physical qubits and a logical error rate per round of $\epsilon_{L}=1.43\times 10^{-4}$, we find a $[[144,12,12]]$ BB code using 576 qubits that matches the performance, with $\epsilon_{L}=1.56\times 10^{-4}$. Additionally, we find a $[[196,12,8]]$ code using 784 qubits that outperforms it with $\epsilon_{L}=7.89\times 10^{-5}$. At this scale, the improvements are not so drastic, but we expect to see greater overhead benefits as the block length and number of logical qubits increase.
We now vary the interval at which the long-range generators are measured, the results of which are also shown in Fig. 11(c). For the $[[90,8,6]]$ code there are 44 (not necessarily independent) generators of a single type. Using a routing schedule that was found by hand, all 35 short-range generators of a single type can be routed, purified, and measured in 9 steps; whereas it takes 5 steps to route, purify, and measure the remaining 10 long-range generators of the same type. Measuring the 35 short-range and 10 long-range generators of the opposite type requires an additional 9 and 5 steps, respectively. This means that measuring the long-range generators every five error correction rounds requires a circuit depth that is 28.5% shorter than if the long-range generators were measured every round ($4\cdot 2\cdot 9+2\cdot(5+9)=100$ steps versus $5\cdot 2\cdot(5+9)=140$ steps), at a negligible increase in the logical error rate per round from $\epsilon_{L}=8.41\times 10^{-4}$ to $8.92\times 10^{-4}$. Increasing the size of the code to $[[196,12,8]]$, we again see negligible differences in logical error per round performance between the two measurement schedules, from $\epsilon_{L}=7.41\times 10^{-5}$ to $7.89\times 10^{-5}$, with the additional benefit of a 32.0% decrease in the depth of the syndrome extraction circuit when measured every five rounds. The two codes consist of $22.22\%$ and $35.71\%$ long-range generators, respectively, yet both remain consistent between long-range measurement intervals. As the block length increases, so does the discrepancy between the measurement times of the short- and long-range generators, increasing the circuit depth savings. Additionally, this discrepancy would disproportionally introduce more errors during the long-range measurement rounds, potentially making it more performant to measure these large generators even less frequently. However, this behavior is very dependent on the idle error rate, and changing the idle error rate may cause the two curves to deviate.
@@
Achieving more significant reductions in circuit depth requires measuring the long-range generators much less frequently, as shown in Fig. 8. We display the potential circuit depth savings for two BB code instances as a function of how many error correction rounds elapse between measurements of the long-range generators. The horizontal red lines indicate the maximum potential savings, corresponding to a schedule where the long-range generators are only measured once at the end of the circuit. For example, the $[[144,12,12]]$ BB code requires 16 (15) steps to measure the short- (long)-range generators of a single type, which gives a maximum depth savings of 48.4%. When measuring every five rounds, as in Fig. 11(a)–(b) and Fig. 12, we see a circuit depth savings of 38.7%, indicated by the vertical gray line. Measuring the long-range generators very infrequently will significantly degrade the error correction performance and may not be worth the reduced circuit depth. Instead, it may be more advantageous to measure the long-range generators relatively frequently, e.g., every 2–5 rounds; in that regime we still see considerable circuit depth savings (50% to 80% of the theoretical maximum), but the impact on the logical error rate is negligible.
$[[n,k,d]]$ Qubits $\epsilon_{L}$
$\mathbf{[[72,8,6]]}$ 288 $3.6\times 10^{-4}\pm 4.9\times 10^{-6}$
$\mathbf{[[90,8,6]]}$ 360 $2.0\times 10^{-4}\pm 3.3\times 10^{-6}$
$[[200,8,5]]$ 392 $2.0\times 10^{-4}\pm 6.5\times 10^{-7}$
$\mathbf{[[120,8,8]]}$ 480 $1.3\times 10^{-5}\pm 8.2\times 10^{-7}$
$[[288,8,6]]$ 568 $9.5\times 10^{-5}\pm 2.5\times 10^{-7}$
$\mathbf{[[150,8,8]]}$ 600 $5.9\times 10^{-6}\pm 2.4\times 10^{-7}$
$[[392,8,7]]$ 776 $2.0\times 10^{-5}\pm 1.5\times 10^{-7}$
$[[512,8,8]]$ 1016 $7.5\times 10^{-6}\pm 4.3\times 10^{-8}$
Table 3: Code parameters, total number of qubits used, and $\epsilon_{L}$ as extracted from Eq. (9) for the simulations described in Section IV.4 albeit with no idle error. Code parameters shown in bold correspond to BB code instances. Code parameters not in bold correspond to copies of the rotated surface code.
@@
Even with the reduced idle error rate that we consider here, idle errors are a significant error source, especially on rounds where the long-range generators are measured. In Table 3 and Fig. 12, we perform the same simulations as described above but do not apply idling errors. Due to their short syndrome extraction circuit depths, the surface codes are unaffected from the decrease in idle error. However, we now find that all BB code instances achieve better logical error rates than surface codes while using fewer physical qubits, as shown in Fig. 7. Indeed, the $[[150,8,8]]$ code using 600 physical qubits sees an $\times8$ improvement in the logical error rate per round, from $\epsilon_{L}=5.31\times 10^{-5}$ to $5.9\times 10^{-6}$, and now outperforms eight patches of a $[[64,1,8]]$ rotated surface code using $1016$ physical qubits with $\epsilon_{L}=7.5\times 10^{-6}.$ Achieving negligible idle error rates may not be feasible, but it illustrates the regime where our protocol performs best.
@@ V Discussion
We do find BB codes where the same parity check structure is shared between codes of increasing block length. For code families with this property, the number of steps in the syndrome extraction circuit is constant, so the noise per syndrome extraction cycle coming from idle error does not increase. However, this also means that the percentage of long-range generators and, by extension, the amount of nonlocality in the code, decreases. References [13, 14] showed that it is impossible to beat the asymptotic scaling of the surface code parameters without increasing the amount of nonlocality. Increasing the block length will yield larger $k$ and $d$, but the asymptotic scaling of these codes will approach that of the surface code; however, for finite sizes we would still expect to see significant space overhead savings compared to alternative, lower-rate codes. Even for BB codes with increasing generator shape, it is feasible that the increased error correction capabilities will outpace the increase in idle error. In particular, Fig. 3 shows a power-law relationship between block length and routing depth. Assuming the code is operating below threshold, the exponential suppression in logical error should be sufficient to handle the increased effective idle error.
@@
If the aforementioned issues can be solved, then scaling up should increase the advantage of qLDPC codes over the surface code. One potential solution is to improve the circuit depth of the protocol. An architectural feature that could accomplish this is the ability to perform two-qubit gates on qubits that are some distance $R$ apart [19]. This is a natural operation on neutral-atom devices, where Rydberg-Rydberg interactions, especially dipolar ones [89], can be quite long-range. Furthermore, Rydberg-Rydberg interactions can help with syndrome extraction by naturally realizing long-range generalized multi-control multi-target CNOT gates [90]. At the same time, such long-range Rydberg-based gates may harm parallelism since only one such gate can be implemented within the Rydberg blockade radius at a time. Gates beyond nearest-neighbor could also be feasible in superconducting devices through the use of medium-range couplers or photonic interconnects [91]. When $R$ is a constant, the asymptotic behavior will remain unchanged; however, practically this would mean that the short-range generators would be much easier to implement. With an appropriate choice of $R$, it would then be possible to use the depth-7 circuit of Ref. [18] to measure the short-range generators, in which case the only difficulty would be to measure the long-range generators in the proposed manner. An alternative approach would be to add additional ancilla layers to the architecture. Although this would further increase the qubit overhead, it would allow for more parallelization during the syndrome measurement, decreasing the total circuit depth.
@@ Appendix A Bivariate bicycle code embeddings
Figure 9: Example five-step schedule to route and purify the Bell pairs needed to measure the short-range, $Z$-type generators of a $[[36,4,4]]$ BB code constructed with $\ell=6,m=3$ and by matrices $A=x+y^{3}+y^{2},B=y^{3}+x^{5}+x^{4}$. The first panel shows the structure of the $Z$-type checks (yellow squares) and $X$-type checks (blue squares), as outlined in gray.
Figure 10: Illustration of the long- and short-range generators of a single type for the $[[120,8,8]]$ BB code constructed with $\ell=12,m=5$ and by matrices $A=x^{10}+y^{4}+y,B=1+x+x^{2}$. The qubits contained in the check are highlighted in gray. The checks (blue squares) highlighted in red are the long-range checks, as they cross the long boundary. For both $Z$- and $X$-type checks, there are 12 long-range checks out of a total of 48, yielding a mask percent of 25%.